OpenStudy (anonymous):

Hi I'm working on a probability problem that involves 3 parts. I know how to do part a and b, but I'm stuck on part c. I was wondering if I can get some help. I'll post the question and the attempt for the solution of part c at the bottom. Any help will be great. Thanks! Will give medal.

OpenStudy (anonymous):

Let X=Y-Z where Y and Z are nonnegative random variables such that YZ=0. (a) show that\[cov(Y,Z)\le0\] (b)show that \[var(X)\ge var(X)+var(Z) \] (c) Using the result of part (b) show that \[var(X)\ge var(\max\{0,X\})+var(\max\{0,-X\})\]

OpenStudy (anonymous):

We never learned variance of the maximum of two random variables in my class. After reading online I found something that stated \[var(\max X_i)\le \sum_{i} var(X_i)\] for iid variables. However I couldn't find any other formula. So my attempt for part (c) was to show that \[var(\max\{0,X\})+var(\max\{0,-X\})\le var(0)+var(X)+var(-X)+var(0)\]\[=var(X)+var(-X)=var(X)+var(X)= 2var(X)\] However this goes against what I'm trying to show. I can't seem to find another formula to implement that will allow me to show that. Am I on the right track or far off? Thank you!

OpenStudy (phi):

does YZ=0. mean when Y is >0 then Z=0 and vice versa? also, is this true \[var(X)\ge var(X)+var(Z) \] aren't variances always positive?

OpenStudy (anonymous):

Yes one of them can be zero or both because they are nonnegative Sorry I made a mistake it was supposed to be var(x)>=Var(y)+var(z)

OpenStudy (anonymous):

I'm not sure if they are always positive. I'll try to check my book

OpenStudy (phi):

var = std^2

OpenStudy (phi):

so 0 or positive

OpenStudy (anonymous):

I see. However, I'm not sure how the variance always being positive helps?

OpenStudy (phi):

X is non-negative, so \( 0 \le X\) the max of 0 and X should be X ? i.e. max(0,X)= X so var(max(0,X))= var(X)

OpenStudy (anonymous):

So for var(max(0,-x)) it would be var(0)? Would that give the it is less than or equal to var(x)+var(0) which is var(x)?

OpenStudy (anonymous):

Since x is nonnegative 0 is greater than -x?

OpenStudy (anonymous):

But if Z is greater than y wouldn't x be negative?

OpenStudy (phi):

var(0) is 0 (i.e. the variance of a string of constants)

OpenStudy (anonymous):

var(x)+var(0)= var(x) However is there a possibility that var(max(0,-x))=var(-x)?

OpenStudy (phi):

max(0,-x) returns the max of 0 or -x x is 0 or +, so -x is 0 or - if x is 0 the max of (0,0) is 0 if x is neg, the max of (0, neg) is 0 so you always get 0

OpenStudy (phi):

are sure you don't have a typo in the part c? it would make more sense if they reference Y or Z (and not X) on the right-hand side.

OpenStudy (anonymous):

Why wouldn't it be -(-value of x) meaning -x is the max? No typo. That also through me off because I don't see how to incorporate them.

OpenStudy (phi):

I should have said: -x is 0 or neg max(0,-x) will always be max(0,0) or max(0, neg number) in either case you get 0

OpenStudy (anonymous):

May I ask how you know that x should be nonnegative? Since the problem states that y and z are nonnegative how does that imply x is also nonnegative?

OpenStudy (phi):

oh, I misread that as all were non-negative.

OpenStudy (anonymous):

Does that change the problem?

OpenStudy (anonymous):

Do I have to do more analysis on Y and Z?

OpenStudy (phi):

yes, it looks like we should (at least it makes more sense now)

OpenStudy (anonymous):

So if both y and z are 0 than the max is var(0) for both? If Y greater than Z then Z must be 0 and X is nonnegative meaning Var(x)+var(0)=var(x) and if z is greater that means var(0)+var(-x)=var(x)?

OpenStudy (anonymous):

meaning it can equal 0 or var(x)

OpenStudy (anonymous):

so var(x) is greater than or equal to? Or is my analysis wrong?

OpenStudy (phi):

I have to think about it, but it seems max(0,x) means max(0,y) because y-z is y or -z. y is + , -z is neg,

OpenStudy (phi):

and max(0,y) is y

OpenStudy (anonymous):

That makes sense to me. But when I try to do max(0,-x) using your analysis I'm not sure how it works? So y would probably be 0 ands would be -z meaning -x =z?

OpenStudy (phi):

exactly

OpenStudy (anonymous):

Since y is nonnegative. I understand that one.

OpenStudy (anonymous):

So is this all the analysis that would be required?

OpenStudy (anonymous):

Also max(0,z)=z for the same reason

OpenStudy (phi):

yes so we get from v(x) >= v( max(0,x) ) + v( max(0,-x) ) v(x) >= v( max(0,y) ) + v( max(0,z) ) v(x) >= v(y) + v(z) which is proved in part b.

OpenStudy (anonymous):

Thank you so much! I understand it now. I would have never though of how to implement y or z though. Thank you once again:D

OpenStudy (phi):

it took a while, but muddling along sometimes works.

OpenStudy (anonymous):

Yes. I think I would have kept trying what I was doing over and over again and I would have gotten nowhere though

OpenStudy (anonymous):

Thank you and have a nice day!

OpenStudy (phi):

yw